OpenStudy (anonymous):

Lisa will make punch that is 25% fruit juice by adding pure fruit juice to a 2-liter mixture that is 10% pure fruit juice. How many liters of pure fruit juice does she need to add?

OpenStudy (anonymous):

Well, to find the answer to this, first we add the percents of pure fruit juice. so 25% + 10% = 35%. Then to find how much of 2 liters is 35%, first make 35% into a fraction so we can make a proportion. so 35% becomes: 35/100. so since 100% of the punch is 2 liters, 2 would be the denominator and since you're trying to find 35% OF 2 liters, the numerator would be represented as x so the equation would become: 35/100 = x/2 then you cross multiply so the equation becomes 100x = 35 * 2 so then you multiply 35 with 2 so the equation becomes: 100x = 70 then divide 100 on both sides so the answer would become x = 0.7 so 35% of the 2 liters would be 1.7 liters i hope this helped :) if you need more help or further explanations, please ask :)

OpenStudy (anonymous):

sorry but the 1.7 liters is not in the choices..the choices are as follows: .4 liter, .5 liter, 2 liters, 8 liters...clearly 2 and 8 liters are not the answers so I just have to focus how to get either .5 or .4... thnks by the way ^_^

OpenStudy (anonymous):

OH WHOOPS. I DIDN'T MEAN 1.7 MY BAD, I TYPED IT WRONG, I'M SORRY.

OpenStudy (anonymous):

well, now that you tell me the choices, i know i did it wrong, i'm sorry x_x

OpenStudy (anonymous):

well, in the way i did it before, instead of it being 35 over 100 = x over 2, it becomes 25/100 = x/2. so if you cross multiply and solve, it becomes 0.5 as the answer. i'm sorry for my mistake x_x

OpenStudy (anonymous):

you have two liters that is 10% fruit juice so you have 10% of 2 = .2 liters of fruit juice. you add say \(x\) liters of 100% fruit juice so you have \[.2+x\] liters of fruit juice and \(2+x\) liters total that you want to 25% fruit juice. set \[.2+x=.25(2+x)\] and solve for \(x\)

OpenStudy (anonymous):

probably easiest to multiply by 100 to get rid of the annoying decimals and solve \[20+100x=25(2+x)\] \[20+100x=50+25x\] \[75x=30\] \[x=\frac{75}{30}=\frac{5}{2}=2.5\]

OpenStudy (anonymous):

yey! thank you!!! so its 0.4

OpenStudy (anonymous):

i got 0.5 o_o

OpenStudy (anonymous):

actually i think it is 2.5, which is what i wrote

OpenStudy (anonymous):

I actually checked it because I have the answer key, the answer is 0.4 liter....thanks guys!

OpenStudy (anonymous):

ohh...soryr i can't really help :(

OpenStudy (anonymous):

lol i solve upside down!

OpenStudy (anonymous):

\[75x=30\] \[x=\frac{30}{75}=.4\]

OpenStudy (anonymous):

its okay, I actually solved for it using the equation you gave...